OpenStudy (anonymous):

enter the equation of the circle in standard form with center and radius given. Center (7, 5), radius = 4

OpenStudy (mertsj):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mertsj):

(h,k) is the center. r is the radius

OpenStudy (anonymous):

x^2+y^2-14x-10y+49+25-16=0 x^2+y^2-14x-10y+58=0